2025-11-14 Early Operations Update

Week 3 of the Early Operations period continued efforts to characterize and improve the distribution of delivered image quality.

On-sky testing for the active optics system (AOS) included ongoing validation of the filter focus offset look-up table to evaluate the repeatability of the corrections associated with filter changes. The team also continued a series of stability tests with a fixed optical configuration for each iteration of the test, with the objective to sample variations of the wavefront estimation for a variety of environmental conditions (e.g., telescope pointing with respect to wind direction) as well as telescope elevation and azimuth, and Camera physical rotator angle. Additional on-sky tests were aimed at characterizing the AOS closed loop performance, including running the AOS closed loop for many iterations at fixed location, and as a separate test, intentionally moving the system away from a good optical state and evaluating the speed and repeatability of convergence.

Two more dome louvers were qualified for nighttime operation, such that four total are now available. As the number of active dome louvers is increased, the team is looking ahead at how to optimize the operation of the louvers during the night for stable delivered image quality. The team continued efforts to characterize thermal sources near the optical path during nighttime operation, e.g., acquiring thermal infrared photographs of the top end assembly and other telescope components during the night.

The team continued using the Feature Based Scheduler (FBS) with the LSST scheduler configuration on most nights, typically for 1-2 hours, to provide regular monitoring of the overall system performance. During recent weeks, the typical estimated system contribution to the delivered image quality has been approximately 0.5 arcseconds.

As we enter dark time, the y filter was swapped for u on 11 November such that the currently installed filter set is ugriz.

1 Like